OpenStudy (anonymous):

find the solution of the equation: 2ln(2+x)=1

OpenStudy (anonymous):

\[2\ln (x+1)=1\]\[\ln (2+x)=1/2\] \[e ^{\ln (2+x)}=e ^{(1/2)}\] \[2+x=e ^{(1/2)}\]\[x=(e ^{(1/2)})/2\]

OpenStudy (anonymous):

Actually, x = e^(1/2) - 2 = -.35

OpenStudy (anonymous):

Should be \[x=e ^{(1/2)}-2\]

OpenStudy (anonymous):

I saw it right after I posted.

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

you're welcome